Please help me with these two problems!
Jose has $20,000 to invest. He invests x dollars at 4% and the rest at 4.5%.
Write an equation for the total amount of money Jose will have in one year.
Suppose at the end of the year, Jose has $20,860. How much money did Jose invest in each account?
Thank you sooo much!!!

Answer by stanbon(57361)

Please help me with these two problems!
Jose has $20,000 to invest. He invests x dollars at 4% and the rest at 4.5%.
Write an equation for the total amount of money Jose will have in one year.
Suppose at the end of the year, Jose has $20,860. How much money did Jose invest in each account?
-----------------------
Equations:
Quantity Eq.: x + y = 20,000
Interest Eq.:0.040x+0.045y = 860
--------------------------------
Multiply thru Quantity by 40
Multiply thru Interest by 1000
----------
40x + 40y = 800,000
40x+45y = 860,000
--------------------------------
Subtract and solve for "y":
5y = 60,000
y = 12,000 (amt. invested at 4.5%
---
Solve for "x":
x + y = 20,000
x = 28,000 (amt. invested at 4%)
============================
Cheers,
Stan H.
===========================